40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15610     Entry time: Sun Oct 4 15:32:21 2020
Author: gautam 
Type: Update 
Category: SUS 
Subject: Suspension health check 

Summary:

After the earthquake on September 19 2020, it looks to me like the only lasting damage to suspensions in vacuum is the ETMY UR magnet being knocked off. 

Suspension ringdown tests:

I did the usual suspension kicking/ringdown test:

  • One difference is that I now kick the suspension "N" times where N is the number of PSD averages desired. 
  • After kicking the suspension, it is allowed to ring down with the damping disabled, for ~1100 seconds so that we can get spectra with 1mHz resolution.
  • We may want to get more e-folding times in, but since the Qs of the modes are a few hundred, I figured this is long enough.
  • I think this kind of approach gives better SNR than letting it ringdown 10,000 seconds (for 10 averages with 10 non overlapping segments of 1000 seconds), and I wanted to test this scheme out, seems to work well.
  • Attachment #1 shows a summary of the results.
  • Attachment #2 has more plots (e.g. transfer function from UL to all other coils), in case anyone is interested in more forensics. The data files are large but if anyone is interested in the times that the suspension was kicked, you can extract it from here.

Conclusions:

  1. My cursory scans of the analysis don't throw up any red flags (apart from the known problem of ETMY UR being dislodged) 👌 
  2. The PRM data is weird 
    • I believe this is because the DC bias voltage to the coils was significantly off from what it normally is when the PRC is aligned.
    • In any case, I am able to lock the PRC, so I think the PRM magnets are fine.
  3. The PRC angular FF no longer works turns out this was just a weird interaction with the Oplev loop because the beam was significantly off-centered on the Oplev QPD. Better alignment fixed it, the FF works as it did before.
    • With the PRC locked and the carrier resonant (no ETMs), the old feedforward filters significantly degrade the angular stability to the point that the lock is lost.
    • My best hypothesis is that the earthquake caused a spot shift on PR2/PR3, which changed the TF from seismometer signal to PRC spot motion.
    • Anyways, we can retrain the filter.
    • The fact that the PRC can be locked suggest PR2/PR3 are still suspended and okay.
  4. The SRM data is also questionable, because the DC bias voltage wasn't set to the values for an aligned SRC when the data was collected
    • Nevertheless, the time series shows a clean ringdown, so at least all 5 OSEMs are seeing a signal.
    • Fact that the beam comes out at the AS port suggest SR3/SR2 suspensions are fine 👍 

Attachment #2 also includes info about the matrix diagonalization, and the condition numbers of the resulting matrices are as large as ~30 for some suspensions, but I think this isn't a new feature. 

Attachment 1: combined.pdf  3.195 MB  Uploaded Sun Oct 4 16:38:03 2020  | Hide | Hide all
combined.pdf
Attachment 2: allPlots.zip  10.274 MB  Uploaded Sun Oct 4 16:50:04 2020
ELOG V3.1.3-